he cost of equipment purchased by Charleston, Inc., on June 1, 2014, is $194,020. It is estimated that the machine will have a $10,900 salvage value at the end of its service life. Its service life is estimated at 7 years, its total working hours are estimated at 91,560, and its total production is estimated at 1,144,500 units. During 2014, the machine was operated 13,080 hours and produced 119,900 units. During 2015, the machine was operated 11,990 hours and produced 104,640 units.

Compute depreciation expense on the machine for the year ending December 31, 2014, and the year ending December 31, 2015, using the following methods.

(a)

Straight-line for 2014 and 2015

(b) Units-of-output for 2014 and 2015
(c) Working hours for 2014 and 2015
(d)

Sum-of-the-years'-digits for 2014 and 2015

(e) Double-declining-balance (twice the straight-line rate) for 2014 and 2015
Answer #1

a . Depriciation as per Straight line method = cost - salvage vaue / life

For 2014 = (194020-10900/7)x7/12 = $15260.

For 2015 = (194020-10900/7) = $26160.

b. Depriciation as per units of output method = (cost - salvage value) x actual production during the period / estimated production during life time

For 2014 = (194020-10900)x119900/1144500 = $19184.

For 2015 = (194020-10900)x104640/1144500 = $16742.4

c . Depriciation as per working hours method = (cost - salvage value )xactusl working hours for the period / total working hours during life time

For 2014 = (194020-10900)x13080/91560

= $26160.

For 2015 = (194020-10900)x11990/91560 = $23980.

d. Depriciation as per sum of digit method

In the given case sum of years = 1+2+3+4+5+6+7 = 28

Depriciation for 2014 = (194020-10900)x7/28x7/12 = $26705

Depriciation for 2015 = (194020-10900)x7/28x5/12+(194020-10900)x6/28x7/12 = $41965

e . Depriciation as per double declining method

Double declining depriciation rate = (1/7)x2 = 28.57%.

Depriciation for 2014 = 194020x28.57%x7/12 =$32335.

Depriciation for 2015 = (194020-32335) x28.57% = $46193.

Note :- in case of double declining balance method there is no need to reduce salvage value.